OpenStudy (anonymous):

The solid described is similar to a larger solid with the given scale factor. Find the surface area and volume of the larger solid A right cylinder with a surface area of 48pi square centimeters and a volume of 45pi cubic centimeter; Scale factor 2:3

OpenStudy (anonymous):

find the radius and height of the cylinder using the surface area and volume provided. since the scale is 2:3, you have to multiply the radius and height by 3/2 and to get the radius and height of the bigger cylinder. Use those values to calculate volume and S.A of the bigger cylinder.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

what didn't you understand in the explanation I gave?

OpenStudy (anonymous):

none of it.. sorry ive been doing math all day

OpenStudy (anonymous):

surface area of the given cylinder is 48pi. so 48 pi = 2pi * radius^2+ 2pi radius*height volume is 45pi so 45 = pi*radius^2*height solve the two equations to get radius and height new radius = 3/2 * radius new height = 3/2 * height use the new values to find the volume and surface area of the new bigger cylinder.

OpenStudy (anonymous):

sorry it is 45pi = pi*radius^2*height